OpenStudy (anonymous):

A ladder is resting against a wall. The ladder and the ground make an angle of 35° and the ladder is 8 ft from the wall. How long is the ladder? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

SOHCAHTOA Do you know these three equations? \[sin \theta = \frac{opposite}{hypotenuse}\]\[cos\theta = \frac{adjacent}{hypotenuse}\]\[tan\theta = \frac{opposite}{adjacent}\]